OpenStudy (anonymous):

What are the inverse operations necessary to solve the equation four times s minus nine equals thirteen?

OpenStudy (anonymous):

4s - 9 = 13

OpenStudy (anonymous):

Add nine to both sides in order to retain terms with s on the left side without any terms with s on the righthand side. After that, you divide by 4 to end up with something that looks like s = ... Problem solved. Is this what you are looking for ?

OpenStudy (uri):

@iceicebaby Add 9 both sides 4s - 9 = 13 +9 +9 ========= 4s=22 s=22/4 divide 22 with 4 s=?

OpenStudy (anonymous):

thank you uri

OpenStudy (uri):

Yw :)